### v1

# Incompatibility of rank-one POVMs and uncertainty relations

Abstract.
The incompatibility between quantum measurements is a fundamental feature of quantum mechanics with profound implications for uncertainty relations and quantum information processing. In this paper, we realized that the $s$-order incompatibility of measurements by Xu (inspired by De Bièvre’s “complete incompatibility” ), can be equivalently characterized by spark of the associated measurement matrix. This provides a new linear algebraic perspective on measurement incompatibility. Leveraging this insight, we extend $s$-order incompatibility to rank-one POVMs and establish novel uncertainty relations. Furthermore, we explore the incompatibility of multiple bases and its connection to support uncertainty relations. These results may have applications in quantum cryptography, quantum state reconstruction, and quantum compressed sensing.
1. Introduction
The incompatibility of measurements is a fundamental feature of quantum mechanics, distinguishing quantum phenomena from classical physics. Recent efforts to rigorously quantify incompatibility have revealed profound connections to uncertainty relations, quantum state reconstruction, and the underlying structure of quantum computational advantage.
The notion of ‘incompatibility’ has different formulations in literature, see e.g. the survey paper [8]. Traditionally, incompatibility has been analyzed through the commutator of two quantum operators. However, recent studies have highlighted the significance of the anticommutator or Jordan product as an equally important descriptor of measurement incompatibility. This shift in perspective is motivated by the mathematical fact: for non-negative rank-one operators $A$ and $B$, the non-negativity condition $\{A,B\}\geq 0$ holds if and only if they commute, i.e., $[A,B]=0$. This equivalence reveals a deep duality between algebraic noncommutativity and geometric negativity in operator space – a connection with far-reaching implications for characterizing quantum coherence, see e.g. [5] and [10].
In a different approach, De Bièvre in [1] introduced the concept of complete incompatibility for two orthonormal bases of a Hilbert space $\mathscr{H}$, linking it to the support uncertainty inequality and the identification of non-classical quantum states whose Kirkwood–Dirac quasiprobability distributions fail to be proper probability distributions. Such states have been shown to provides advantages in quantum metrology and have important implications in weak measurements and contextuality. Xu subsequently generalized in [4] this concept to $s$-order incompatibility for a positive integer $s$, where the case $s=d+1$ recovers De Bièvre’s original definition. A key distinction of $s$-order incompatibility is that it takes only a finite set of integer values, contrasting with conventional continuous measures of incompatibility, such as those concepts based on Jordan product in [5].
We observe that the complete incompatibility or the more general $s$-order incompatibility for two orthonormal bases can be characterized using a concept from linear algebra — the spark of the matrix formed by concatenating the basis vectors. In the context of compressed sensing, the spark represents the smallest number of linearly dependent columns in a matrix, providing a concrete and computable measure of incompatibility. This insight offers a novel linear algebraic perspective on measurement incompatibility, linking it to sparsity and compressed sensing techniques.
Extending $s$-order incompatibility to general positive operator-valued measures (POVMs) via tight frame theory [12], we address the challenge of characterizing non-projective measurement incompatibility. This extension reveals that the incompatibility of POVMs is governed by the linear dependence structure of their measurement vectors, making it particularly useful where non-projective measurements are essential, such as quantum state estimation with noise resilience. Consequently, we derive a new class of uncertainty relations that emphasize the interplay between measurement incompatibility and the minimal support of quantum states across different measurement settings. Notably, we offer a novel interpretation of the Ghobber-Jaming uncertainty relation [3] through the lens of measurement incompatibility. This interpretation is further generalized to rank-one POVMs, addressing a question raised in [15] for general frames.
Furthermore, we extend our framework to systems involving multiple orthonormal bases, introducing a generalized $s$-order incompatibility criterion. Recall that variance-based uncertainty relations have been established for multiple observables associated with different bases, see e.g. [13]. Our approach enables the formulation of sharp multi-measurement support uncertainty relations. Such relations are crucial in applications requiring compatibility constraints across several observables, such as quantum error correction and multi-parameter metrology. Unlike the two-bases case, where the spark serves as a useful tool, the multi-bases setting requires a different generalization of the concepts by De Bièvre and Xu. Notably, our support uncertainty relation has sharp lower bound, in contrast to the continuous setting, where establishing sharp bounds for multiple measurements remains an open challenge.
These findings not only deepen our understanding of measurement incompatibility but also have potential applications in quantum cryptography, quantum state reconstruction, and quantum error correction. In particular, the spark-based characterization may provide new insights into designing and analyzing measurement schemes in quantum compressed sensing and quantum tomography.
Paper Organization and Notations: The remainder of this paper is structured as follows. In Section 2 establishes the equivalence between $s$-order incompatibility and the spark of measurement matrix. Section 3 extends this concept to POVMs via frame theory. Section 4 derives novel uncertainty relations within the incompatibility framework. Section 5 explores the implications of incompatibility in multi-basis settings and its connection to support uncertainty. Finally, we conclude by discussing potential applications and open questions.
Throughout the paper, the set of consecutive integers $\{1,2,\ldots,d\}$ is denoted by $[[1,d]]$. For two orthogonal bases $A$ and $B$, we denote $(A,B)$ by the matrix whose columns are given by the basis of $A$ and $B$. For a set $S$, the notion $|S|$ denotes the number of elements in $S$, ${\rm span}\{S\}$ denotes the space spanned by $S$ over the complex field $\mathbb{C}$. Let $I=\{1,2,\ldots,m\}$ and $J=\{1,2,\ldots,n\}$. For a set $E\subset I$, we will write $E^{c}$ for its complement.
2. $s$-order incompatibility and the spark of Measurement Matrices
Let $A=\{\ket{a_{j}}\}_{j=1}^{d}$ and $B=\{\ket{b_{k}}\}_{k=1}^{d}$ be two orthonormal bases for a $d$-dimensional complex Hilbert space $\mathscr{H}$. As usual to avoid the freedom |$\ket{a_{j}}\rightarrow e^{i\theta_{j}}\ket{a_{j}}$ with $\theta_{j}\in\mathbb{R}$ real number and $i=\sqrt{-1}$, we consider the sets of rank-one projectors $\overline{A}=\{\ket{a_{j}}\bra{a_{j}}\}_{j=1}^{d}$ and $\overline{B}=\{\ket{b_{k}}\bra{b_{k}}\}_{k=1}^{d}$. When $\overline{A}$ and $\overline{B}$ commute, we say $A$ and $B$ are compatible. Otherwise we say $A$ and $B$ is incompatible. De Bièvre introduced the concept of complete incompatibility in [1], which is purely algebraic but has a clear physical interpretation in terms of selective projective measurements. It has been shown that this concept closely related to minimal support uncertainty, and it also is useful to characterize the Kirkwood-Dirac nonclassicality. The incompatibility of two bases is further classified by Xu in [4] which includes De Bièvre’s complete incompatibility as a special case.
Definition 2.1 ($s$-order incompatibility [4]).
Two orthonormal bases $A$ and $B$ are $s$-order incompatible, where $s\in[[2,d+1]]$, if
  1. (1)
    For all nonempty subsets $S_{A}\subset A$ and $S_{B}\subset B$ with $|S_{A}|+|S_{B}|<s$, then
    ${\rm span}\{S_{A}\}\cap{\rm span}\{S_{B}\}=\{0\}.$
  2. (2)
    There exist subsets $S_{A}$ and $S_{B}$ with $|S_{A}|+|S_{B}|=s$, such that
    ${\rm span}\{S_{A}\}\cap{\rm span}\{S_{B}\}\neq\{0\}.$
On the other hand, we recall the spark of a matrix from compressed sensing in [11] by D.L. Donoho and M. Elad.
Definition 2.2 (Spark or Kruskal rank).
The spark of a $m\times n$ matrix $M$ is the smallest integer $k$ such that there exists a set of $k$ columns in $M$ which are linearly dependent.
Now, we have the following observation. The proof is omitted.
Theorem 2.3.
Two orthonormal bases $A$ and $B$ are $s$-order incompatible if and only if ${\rm spark}(A,B)=s,$ where $(A,B)$ is the matrix whose column vectors are the basis vectors.
3. Incompatibility of rank-one POVMs via tight frames
A Positive Operator-Valued Measure (POVM) is a set of non-negative Hermitian operators $\{Q_{j}\}$, not necessarily projectors, that satisfy $\sum_{j}Q_{j}={\rm\mathbb{I}}$ where ${\rm\mathbb{I}}$ is the identity operator on $\mathscr{H}$. A POVM generalizes the standard measurement, as its measurement vectors $\mu_{j}$ need not to be normalized or orthogonal. There exists a fundamental relationship between rank-one quantum measurements and the so-called tight frames [12]. Frames generalize bases and were introduced in the context of nonharmonic Fourier series by Duffin and Schaeffer.
Definition 3.1.
Let $\{\varphi_{j},1\leq j\leq n\}$ denote a set of (overcomplete) vectors in a $d$-dimensional space $\mathscr{H}$. The vectors $\{\varphi_{j}\}$ is said to form a tight frame if there exists a constant $\alpha>0$ such that for all $x\in\mathscr{H}$,
$\sum_{j=1}^{n}|\langle x,\varphi_{j}\rangle|^{2}=\alpha^{2}\|x\|^{2}.$
If $\alpha=1$, the tight frame is said to be normalized, otherwise, it is $\alpha$-scaled.
Theorem 3.2.
[12] A set of vectors $\varphi_{j}\in\mathscr{H}$ forms a $\alpha$-scaled tight frame for $\mathscr{H}$ if and only if the scaled vectors $\alpha^{-1}\varphi_{j}$ are the measurement vectors of a rank-one POVM on $\mathscr{H}$. In particular, the vectors $\varphi_{j}$ form a normalized tight frame for $\mathscr{H}$ if and only if they are the measurement vectors of a rank-one POVM on $\mathscr{H}$.
When no confusion can arise, we will use the same notation for the POVMs and the corresponding set of measurement vectors.
Definition 3.3 ($s$-order incompatibility for POVMs).
Two rank-one POVMs $A$ and $B$ are called $s$-order incompatible if ${\rm spark}(A,B)=s$, where $(A,B)$ are the measurement vector matirx.
For a pure state |$\ket{\varphi}$, let $\{\langle a_{j}|\varphi\rangle\}$ and $\{\langle b_{k}|\varphi\rangle\}$ denote the coefficients under the POVMs. Denote $n_{A}(\ket{\varphi})$ for the number of non-zero elements in $\{\langle a_{j}|\varphi\rangle\}$ and $n_{B}(\ket{\varphi})$ as the number of non-zero elements in $\{\langle b_{k}|\varphi\rangle\}$. Then the following sharp support uncertainty relation holds.
Theorem 3.4 (support uncertainty relation).
For $s$-order incompatible rank-one POVMs $A$ and $B$, Then for any non-zero pure state |$\ket{\varphi}$, we have
$n_{A}(\ket{\varphi})+n_{B}(\ket{\varphi})\geq s.$
Proof.
It follows from the property of spark, as detailed in [11]. ∎
This yields an equivalent definition of the $s$-order incompatibility for a pair of POVMs or for general tight frames.
Definition 3.5.
Suppose $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ are two tight frames in $\mathscr{H}$, with frame constant $\alpha$ and $\beta$. If the integer $s$ satisfies
  1. (1)
    For any non-empty subsets $S\subseteq I$ and $J\subseteq J$, satisfying $|S|+|T|<s$, such that for any non-zero $x\in\mathscr{H}$,
    $\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2},\quad{\rm and}\quad\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}$
    do not hold simultaneously,
  2. (2)
    There exist non-empty subsets $S\subseteq I$ and $T\subseteq J$, satisfying $|S|+|T|=s$, such that there exists a non-zero $x\in\mathscr{H}$, the following
    $\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2},\quad{\rm and}\quad\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}$
    hold simultaneously.
Then $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ are called $s$-order incompatible.
4. Uncertainty relations from $s$-order incompatibility
In this section, we establish new uncertainty relation based on the concept of s-order incompatibility. By leveraging the spark of measurement matrices, we derive lower bounds on the support of quantum states under different measurement settings, extending traditional uncertainty principles to a broader class of quantum measurements. Let $\{a_{k}\}_{k=1}^{m}$ and $\{b_{j}\}_{j=1}^{n}$ be two frames in a Hilbert space $\mathscr{H}$, where each set corresponds to a rank-one positive operator-valued measure (POVM).
Definition 4.1 (Minimal Reconstruction Number).
For frames $A$ and $B$, define
$t_{\min}=\min\{t\,|\,\forall S\subsetneq I,T\subsetneq J,|S|+|T|>t,\operatorname{span}\left(\{a_{k}\}_{k\in S}\cup\{b_{j}\}_{j\in T}\right)=\mathscr{H}\}.$
For any pair of $s$-order incompatible tight frames, the relationship between $s$ and $t_{\min}$ is given by the following theorem:
Theorem 4.2.
If two tight frames $A$ and $B$ are $s$-order incompatible, then
$s+t_{\min}=m+n.$
Proof.
We show that $t_{\min}=m+n-s$.
1) For any subsets $S\subsetneq I$ and $T\subsetneq J$ satisfying $|S|+|T|>m+n-s$, it follows that their complements $S^{c}$ and $T^{c}$ satisfy
$|S^{c}|+|T^{c}|<s.$
Since the frames $\{a_{k}\}_{k=1}^{m}$ and $\{b_{j}\}_{j=1}^{n}$ are $s$-order incompatible, thus by the equivalent Definition 3.5, for any non-zero vector $x\in\mathscr{H}$, the following two identities
$\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2},\quad\mbox{and}\quad\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}$
cannot simultaneously hold. This implies that
$\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}+\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}<(\alpha+\beta)\|x\|^{2},\qquad\forall\,x\neq 0.$
It is equivalent to say, for any non-zero $x\in\mathscr{H}$,
$\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}+\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}>0,$
thus $\left(\operatorname{span}\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}\right)^{\perp}=\{0\}$. It follows that
$\operatorname{span}\left(\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}\right)=\mathscr{H}.$
Therefore, $t_{\min}\leq m+n-s$.
2) By the definition of $s$-order incompatibility, there exist two non-empty subsets $S\subset I$ and $T\subset J$ satisfying $|S|+|T|=s$, such that for some nonzero $x$, the following hold
$\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2},\quad\mbox{and}\quad\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}.$
This means that for any $S^{c}\subsetneq I$ and $T^{c}\subsetneq J$ satisfying $|S^{c}|+|T^{c}|=m+n-s$, there exists non-zero $x\in\mathscr{H}$ such that
$\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}+\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}=0.$
Thus $\left({\rm span}\left(\{a_{k}\}_{k\in S}\cup\{b_{j}\}_{j\in T}\right)\right)^{\perp}\neq\{0\}$, i.e.
${\rm span}\left(\{a_{k}\}_{k\in S}\cup\{b_{j}\}_{j\in T}\right)\neq\mathscr{H}.$
Hence $t_{\min}\geq m+n-s$.
Collecting all, we obtain $t_{\min}=m+n-s$, completing the proof. ∎
Remark 4.3.
The quantity $t_{\min}$ establishes a fundamental limit on the number of measurements needed to reconstruct a quantum state from two rank-one POVMs.
Theorem 4.4.
Let $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ in $\mathscr{H}$ be two $s$-order incompatible tight frames. Then there exists a constant $C$ such that
(4.1) $\|x\|^{2}\leq C\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}+\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right).$
for all subsets $S\subseteq I$ and $T\subseteq J$ with $|S|+|T|<s$.
Proof.
By Theorem 4.2, for any non-empty subsets $S\subset I$ and $T\subset J$ satisfying $|S|+|T|<s$, we know that
$\operatorname{span}\left(\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}\right)=\mathscr{H}.$
Thus the set $\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}$ forms a frame for $\mathscr{H}$ (by [14, Lemma 1.2]), whose lower frame bound is denoting by $C_{S,T}$. Denoting $C_{1}=\min C_{S,T}$, the desired constant can be chosen as $C=1/\min\{\alpha,\beta,C_{1}\}$. ∎
Remark 4.5.
If the columns of the matrix $M$ are normalized to unit norm, then the lower bound of its spark can be given by its dictionary coherence
${\rm spark}(M)\geq 1+\frac{1}{\mu(M)},$
where the dictionary coherence $\mu(A)$ is defined as the maximum correlation between any to columns
$\mu(M)=\max_{m\neq n}|\langle a_{m},a_{n}\rangle|.$
It is seen that our uncertainty inequality provides a more essential interpretation for Ghobber-Jaming uncertainty relation in [3] in terms of the coherence $\mu(M)$.
5. Ghobber-Jaming uncertainty relation for frames
Theorem 5.1.
Let $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ be two frames in $\mathscr{H}$, with lower and upper frame bounds $\alpha_{1},\beta_{1}$ and $\alpha_{2},\beta_{2}$, respectively. Suppose that for any subsets $S\subseteq\{1,\dots,m\}$ and $T\subseteq\{1,\dots,n\}$, satisfying $S\subseteq T$, the condition
$\frac{\mu^{2}}{\alpha_{1}}<\frac{\beta_{2}}{T}$
holds, where $\mu(M)=\max_{m\neq n}|\langle a_{m},a_{n}\rangle|$. Then for all $x\in H$,
$C(S,T)\|x\|\leq\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}.$
Proof.
Let $x=x_{1}+x_{2}$, where
$x_{1}=\sum_{k\in S}\langle x,a_{k}\rangle a_{k}^{*},\quad x_{2}=\sum_{k\in S^{c}}\langle x,a_{k}\rangle a_{k}^{*},$
in which $a_{k}^{*}$ is the canonical dual frame.
$\begin{split}\sum_{j\in T}|\langle x_{1},b_{j}\rangle|^{2}=\sum_{j\in T}\left|\sum_{k\in S}\langle x,a_{k}\rangle\langle a_{k}^{*},b_{j}\rangle\right|^{2}\\ \leq\sum_{j\in T}\left(\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}\right)\left(\sum_{k\in S}\left|\langle a_{k}^{*},b_{j}\rangle\right|\right)\\ \leq\mu^{2}\alpha_{2}|T|\|x\|^{2}.\end{split}$
Thus,
$\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}=\sum_{j=1}^{n}\left|\langle x_{1},b_{j}\rangle\right|^{2}-\sum_{j\in T}|\langle x_{1},b_{j}\rangle|^{2}\geq\beta_{1}\|x_{1}\|^{2}-\alpha_{2}\mu^{2}\mu^{2}|T|\|x\|^{2}$
This implies
$\begin{split}\beta_{1}^{\frac{1}{2}}\|x_{1}\|&\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}+\alpha_{2}\mu^{2}|T|\|x\|^{2}\right)^{\frac{1}{2}}\\ &\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\alpha_{2}^{\frac{1}{2}}\mu|T|^{\frac{1}{2}}\|x\|\\ &\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x_{2},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\alpha_{2}^{\frac{1}{2}}\mu|T|^{\frac{1}{2}}\|x\|\\ &\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\beta_{2}^{\frac{1}{2}}\|x_{2}\|+\alpha_{2}^{\frac{1}{2}}\mu|T|^{\frac{1}{2}}\|x\|.\end{split}$
On the other hand,
$\begin{split}\alpha_{1}\|x_{2}\|^{2}&\leq\sum_{k=1}^{m}|\langle x_{2},a_{k}\rangle|^{2}\leq\sum_{k=1}^{m}\left(\sum_{i\in S^{c}}|\langle x,a_{i}\rangle|^{2}\right)\left(\sum_{i\in S^{c}}|\langle a_{i}^{*},a_{k}\rangle|^{2}\right)\\ \leq m^{2}M(A^{*},A)^{2}\sum_{k\in s^{c}}|\langle x,a_{k}\rangle|^{2}\end{split}$
Thus we have
$\begin{split}\|x\|\leq\|x_{1}\|+\|x_{2}\|&\leq\frac{1}{\beta_{1}^{\frac{1}{2}}}\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\left(1+\left(\frac{\beta_{2}}{\beta_{1}}\right)^{1/2}\right)\|x_{2}\|+\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}\mu|T|^{1/2}\|x\|\end{split}$
we obtain
$\begin{split}C(S,T)\|x\|\leq\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{1/2}\end{split}$
where
$C(S,T)=\left(1-\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}\mu|T|^{1/2}\right)\left(\max\{\frac{1}{\beta_{1}^{1/2}},\left\{\beta_{1}^{-\frac{1}{2}},\left((\frac{\beta_{2}}{\alpha_{1}\beta_{1}})^{\frac{1}{2}}+\alpha_{1}^{-\frac{1}{2}}\right)mM(A^{*},A)\right\}\right)^{-1}.$
6. Incompatibility of multiple bases
We consider the incompatibility of multiple bases in this section.
Definition 6.1.
Orthonormal bases $A_{k}=\{\ket{a^{(k)}_{j}}\}_{j=1}^{n}$ are $s$-order incompatible if:
  1. (1)
    For any nonempty $S_{k}\subset A_{k}$ with $\sum_{k=1}^{n}|S_{k}|<s$, $\bigcap_{k=1}^{n}{\rm span}\{S_{k}\}=\{0\}.$
  2. (2)
    There exist subsets $S_{k}\subset A_{k}$, such that $\sum_{k=1}^{n}|S_{k}|=s$ and $\bigcap_{k=1}^{n}{\rm span}\{S_{k}\}\neq\{0\}.$
Theorem 6.2.
For $s$-order incompatible bases $\{A_{k}\}_{k=1}^{n}$, any non-zero pure state |$\ket{\varphi}$ if and only if
$\sum_{k=1}^{n}n_{A_{k}}(\ket{\varphi})\geq s.$
Proof.
Let
$n_{\min}=\min_{\ket{\varphi}\neq 0}\sum_{k=1}^{n}n_{A_{k}}(\ket{\varphi}).$
Then if $n_{\min}=s$, there exists a pure state |$\ket{\psi}$ such that
$\sum_{k=1}^{n}n_{A_{k}}(\ket{\psi})=s,$
and there does not exist a pure state |$\ket{\psi^{\prime}}$ such that $\sum_{k=1}^{n}n_{A_{k}}(\ket{\psi^{\prime}})<s$.
For such |$\ket{\psi}$, there exist nonempty subsets $S_{A_{k}}\subseteq A_{k}$ such that $|S_{A_{k}}|=n_{A_{k}}(\ket{\psi})$, and |$\ket{\psi}\in\bigcap_{k=1}^{n}{\rm span}(S_{A_{k}})$. The nonexistence of such |$\ket{\psi^{\prime}}$ implies that there do not exist nonempty subsets $S_{A_{k}}\subseteq A_{k}$ such that $|S_{A_{k}}|=n_{A_{k}}(\ket{\psi^{\prime}})$, and |$\ket{\psi^{\prime}}\in\bigcap_{k=1}^{n}{\rm span}(S_{A_{k}})$. These two conditions precisely coincide with (1.1) and (1.2) in Definition 1. Then the claim follows. $\square$
7. Conclusion
By equating $s$-order incompatibility with the spark, we unify quantum measurement theory with compressed sensing. Key contributions include:
  • POVM Incompatibility: Generalized to non-projective measurements via tight frames.
  • Sharp Uncertainty Bounds: Derived from the spark, improving coherence-based estimates.
  • Multi-Basis Framework: Extended support uncertainty to multi-observable systems.
Future directions include applications in quantum tomography, cryptographic protocol design, and infinite-dimensional spark generalizations.

### v2

@@
-Incompatibility of rank-one POVMs and uncertainty relations
+Incompatibility of rank-one POVMs and quantum uncertainty relations
 Abstract.
-The incompatibility between quantum measurements is a fundamental feature of quantum mechanics with profound implications for uncertainty relations and quantum information processing.
-In this paper, we realized that the $s$-order incompatibility of measurements by Xu (inspired by De Bièvre’s “complete incompatibility” ), can be equivalently characterized by spark of the associated measurement matrix.
-This provides a new linear algebraic perspective on measurement incompatibility.
-Leveraging this insight, we extend $s$-order incompatibility to rank-one POVMs and establish novel uncertainty relations.
-Furthermore, we explore the incompatibility of multiple bases and its connection to support uncertainty relations.
+The incompatibility of quantum measurements is a fundamental feature of quantum mechanics with profound implications for uncertainty relations and quantum information processing.
+In this paper, we extend the notion of $s$-order incompatibility of measurements, introduced by Xu (inspired by De Bièvre’s “complete incompatibility” ), to more general rank-one POVMs, and establish novel uncertainty relations.
+Furthermore, we investigate the incompatibility of multiple POVMs and its connection to support uncertainty relations.
 These results may have applications in quantum cryptography, quantum state reconstruction, and quantum compressed sensing.
 1. Introduction
-The incompatibility of measurements is a fundamental feature of quantum mechanics, distinguishing quantum phenomena from classical physics.
-Recent efforts to rigorously quantify incompatibility have revealed profound connections to uncertainty relations, quantum state reconstruction, and the underlying structure of quantum computational advantage.
-The notion of ‘incompatibility’ has different formulations in literature, see e.g.
-the survey paper [8].
-Traditionally, incompatibility has been analyzed through the commutator of two quantum operators.
-However, recent studies have highlighted the significance of the anticommutator or Jordan product as an equally important descriptor of measurement incompatibility.
-This shift in perspective is motivated by the mathematical fact: for non-negative rank-one operators $A$ and $B$, the non-negativity condition $\{A,B\}\geq 0$ holds if and only if they commute, i.e., $[A,B]=0$.
-This equivalence reveals a deep duality between algebraic noncommutativity and geometric negativity in operator space – a connection with far-reaching implications for characterizing quantum coherence, see e.g.
-[5] and [10].
-In a different approach, De Bièvre in [1] introduced the concept of complete incompatibility for two orthonormal bases of a Hilbert space $\mathscr{H}$, linking it to the support uncertainty inequality and the identification of non-classical quantum states whose Kirkwood–Dirac quasiprobability distributions fail to be proper probability distributions.
-Such states have been shown to provides advantages in quantum metrology and have important implications in weak measurements and contextuality.
-Xu subsequently generalized in [4] this concept to $s$-order incompatibility for a positive integer $s$, where the case $s=d+1$ recovers De Bièvre’s original definition.
-A key distinction of $s$-order incompatibility is that it takes only a finite set of integer values, contrasting with conventional continuous measures of incompatibility, such as those concepts based on Jordan product in [5].
+The incompatibility of measurements is a fundamental feature of quantum mechanics, setting it apart from classical physics.
+Recent efforts to rigorously quantify incompatibility have revealed profound connections to uncertainty relations, quantum state reconstruction, etc.
+Traditionally, incompatibility has been studied using the commutator of quantum operators [7].
+Recent studies suggest that the alternative operator, i.e.
+anticommutator or Jordan product as an equally important descriptor of measurement incompatibility.
+This can be find some clue from the mathematical fact: for any non-negative rank-one operators $A$ and $B$, the non-negativity condition $\{A,B\}\geq 0$ holds if and only if they commute, i.e., $[A,B]=0$.
+This equivalence highlights a deep interplay between algebraic noncommutativity and geometric negativity in operator space, revealing new perspectives on quantum coherence and measurement incompatibility, see e.g., [6, 8].
+A different approach to characterizing incompatibility was introduced by De Bièvre in [2], who introduced the concept of complete incompatibility for two orthonormal bases of a Hilbert space $\mathscr{H}$, linking it to the support uncertainty inequality and the identification of non-classical quantum states whose Kirkwood–Dirac quasiprobability distributions fail to be proper probability distributions.
+De Bièvre’s notion was further extended by Xu in [16] to the $s$-order incompatibility, which provides a discrete measure (taking only integers numbers) of incompatibility rather than a continuous one, making it significantly different from previous formulations, such as [8].
+Despite these advancements, an interesting question remains: how can we generalize these notions beyond orthonormal bases to more general measurement frameworks such as Positive Operator-Valued Measures (POVMs)?
 We observe that the complete incompatibility or the more general $s$-order incompatibility for two orthonormal bases can be characterized using a concept from linear algebra — the spark of the matrix formed by concatenating the basis vectors.
@@
 This insight offers a novel linear algebraic perspective on measurement incompatibility, linking it to sparsity and compressed sensing techniques.
-Extending $s$-order incompatibility to general positive operator-valued measures (POVMs) via tight frame theory [12], we address the challenge of characterizing non-projective measurement incompatibility.
+The main contribution of this paper is extending the $s$-order incompatibility to general rank-one positive operator-valued measures (POVMs) via frame theory [4], we address the challenge of characterizing non-projective measurement incompatibility.
 This extension reveals that the incompatibility of POVMs is governed by the linear dependence structure of their measurement vectors, making it particularly useful where non-projective measurements are essential, such as quantum state estimation with noise resilience.
-Consequently, we derive a new class of uncertainty relations that emphasize the interplay between measurement incompatibility and the minimal support of quantum states across different measurement settings.
-Notably, we offer a novel interpretation of the Ghobber-Jaming uncertainty relation [3] through the lens of measurement incompatibility.
-This interpretation is further generalized to rank-one POVMs, addressing a question raised in [15] for general frames.
-Furthermore, we extend our framework to systems involving multiple orthonormal bases, introducing a generalized $s$-order incompatibility criterion.
-Recall that variance-based uncertainty relations have been established for multiple observables associated with different bases, see e.g.
-[13].
-Our approach enables the formulation of sharp multi-measurement support uncertainty relations.
-Such relations are crucial in applications requiring compatibility constraints across several observables, such as quantum error correction and multi-parameter metrology.
-Unlike the two-bases case, where the spark serves as a useful tool, the multi-bases setting requires a different generalization of the concepts by De Bièvre and Xu.
-Notably, our support uncertainty relation has sharp lower bound, in contrast to the continuous setting, where establishing sharp bounds for multiple measurements remains an open challenge.
+Building on this framework, we establish a new class of uncertainty relations in Section 4 that emphasize the interplay between measurement incompatibility and the minimal support of quantum states across different measurement settings, in the sprit of the Ghobber-Jaming uncertainty relation [5].
+In Section 5, the original Ghobber-Jaming uncertainty is generalized in a different way using coherence to rank-one POVMs, addressing a question raised in [13] for general frames.
+Furthermore, we extend our framework to systems involving multiple rank-one POVMs, introducing a generalized $s$-order incompatibility criterion.
+Recall that variance-based uncertainty relations have been studied for multiple observables associated with different bases, see e.g., [11, 12].
+Our approach leads to multi-measurement support uncertainty relation, which will be useful in applications requiring compatibility constraints across several observables, such as quantum error correction and multi-parameter metrology.
+Moreover, our support uncertainty relation has sharp lower bound, in contrast to the continuous setting, where establishing sharp bounds for multiple measurements remains open.
 These findings not only deepen our understanding of measurement incompatibility but also have potential applications in quantum cryptography, quantum state reconstruction, and quantum error correction.
-In particular, the spark-based characterization may provide new insights into designing and analyzing measurement schemes in quantum compressed sensing and quantum tomography.
-Paper Organization and Notations: The remainder of this paper is structured as follows.
-In Section 2 establishes the equivalence between $s$-order incompatibility and the spark of measurement matrix.
-Section 3 extends this concept to POVMs via frame theory.
+The spark-based characterization may provide new insights into designing and analyzing measurement schemes in quantum compressed sensing and quantum tomography.
+Paper Organization and Notations: The rest of this paper is structured as follows.
+Section 2 establishes the equivalence between $s$-order incompatibility and the spark of the measurement matrix.
+Section 3 extends the $s$-order incompatibility to rank-one POVMs via frame theory.
 Section 4 derives novel uncertainty relations within the incompatibility framework.
-Section 5 explores the implications of incompatibility in multi-basis settings and its connection to support uncertainty.
-Finally, we conclude by discussing potential applications and open questions.
+Section 5 provides an alternative generalization of Ghobber-Jaming uncertainty relations for frames.
+Section 6 investigate the implications of incompatibility in multi-basis settings and its connection to support uncertainty.
+We conclude this paper by discussing potential applications and open questions.
 Throughout the paper, the set of consecutive integers $\{1,2,\ldots,d\}$ is denoted by $[[1,d]]$.
@@
 For a set $E\subset I$, we will write $E^{c}$ for its complement.
-2. $s$-order incompatibility and the spark of Measurement Matrices
+2. $s$-order incompatibility and spark
 Let $A=\{\ket{a_{j}}\}_{j=1}^{d}$ and $B=\{\ket{b_{k}}\}_{k=1}^{d}$ be two orthonormal bases for a $d$-dimensional complex Hilbert space $\mathscr{H}$.
-As usual to avoid the freedom |$\ket{a_{j}}\rightarrow e^{i\theta_{j}}\ket{a_{j}}$ with $\theta_{j}\in\mathbb{R}$ real number and $i=\sqrt{-1}$, we consider the sets of rank-one projectors $\overline{A}=\{\ket{a_{j}}\bra{a_{j}}\}_{j=1}^{d}$ and $\overline{B}=\{\ket{b_{k}}\bra{b_{k}}\}_{k=1}^{d}$.
-When $\overline{A}$ and $\overline{B}$ commute, we say $A$ and $B$ are compatible.
-Otherwise we say $A$ and $B$ is incompatible.
-De Bièvre introduced the concept of complete incompatibility in [1], which is purely algebraic but has a clear physical interpretation in terms of selective projective measurements.
-It has been shown that this concept closely related to minimal support uncertainty, and it also is useful to characterize the Kirkwood-Dirac nonclassicality.
-The incompatibility of two bases is further classified by Xu in [4] which includes De Bièvre’s complete incompatibility as a special case.
+To eliminate the phase ambiguity |$\ket{a_{j}}\rightarrow e^{i\theta_{j}}\ket{a_{j}}$ with $\theta_{j}\in\mathbb{R}$ real number and $i=\sqrt{-1}$, we consider as usual the sets of rank-one projectors (for a precise definition, see Definition 3.1 below)
+[…]
+The bases $A$ and $B$ are said to be compatible if the projectors $\overline{A}$ and $\overline{B}$ commute; otherwise, they are considered incompatible.
+It is worthwhile to further classify the degrees of incompatibility.
+De Bièvre introduced the purely algebraic notion complete incompatibility in [2], which is further refined by Xu as follows.
 Definition 2.1 ($s$-order incompatibility [16]).
-Two orthonormal bases $A$ and $B$ are $s$-order incompatible, where $s\in[[2,d+1]]$, if
+Two orthonormal bases $A$ and $B$ are said to be $s$-order incompatible, where $s\in[[2,d+1]]$, if
   1. (1)
-    For all nonempty subsets $S_{A}\subset A$ and $S_{B}\subset B$ with $|S_{A}|+|S_{B}|<s$, then
+    For all nonempty subsets $S_{A}\subset A$ and $S_{B}\subset B$ with $|S_{A}|+|S_{B}|<s$, it holds that
     ${\rm span}\{S_{A}\}\cap{\rm span}\{S_{B}\}=\{0\}.$
@@
     ${\rm span}\{S_{A}\}\cap{\rm span}\{S_{B}\}\neq\{0\}.$
-On the other hand, we recall the spark of a matrix from compressed sensing in [11] by D.L. Donoho and M. Elad.
+Remark 2.2.
+This concept provides a characterization for the degree of incompatibility.
+Alternatively, the concept of a matrix’s spark is widely used in compressed sensing, see e.g., [3].
 Definition 2.3 (Spark or Kruskal rank).
-The spark of a $m\times n$ matrix $M$ is the smallest integer $k$ such that there exists a set of $k$ columns in $M$ which are linearly dependent.
-Now, we have the following observation. The proof is omitted.
+The spark of a $m\times n$ matrix $M$ is the smallest integer $k$ such that there exists a set of $k$ columns in $M$ that are linearly dependent.
+It is interesting to observe the following connection.
 Theorem 2.4.
 Two orthonormal bases $A$ and $B$ are $s$-order incompatible if and only if ${\rm spark}(A,B)=s,$ where $(A,B)$ is the matrix whose column vectors are the basis vectors.
-3. Incompatibility of rank-one POVMs via tight frames
-A Positive Operator-Valued Measure (POVM) is a set of non-negative Hermitian operators $\{Q_{j}\}$, not necessarily projectors, that satisfy $\sum_{j}Q_{j}={\rm\mathbb{I}}$ where ${\rm\mathbb{I}}$ is the identity operator on $\mathscr{H}$.
-A POVM generalizes the standard measurement, as its measurement vectors $\mu_{j}$ need not to be normalized or orthogonal.
-There exists a fundamental relationship between rank-one quantum measurements and the so-called tight frames [12].
-Frames generalize bases and were introduced in the context of nonharmonic Fourier series by Duffin and Schaeffer.
+3. Incompatibility of tight frames and rank-one POVMs
+In this section, we classify the incompatibility of more general quantum measurements.
+Recall that a Positive Operator-Valued Measure (POVM) is a set of non-negative Hermitian operators $\{Q_{j}\}$, which are not necessarily projectors, satisfying $\sum_{j}Q_{j}={\rm\mathbb{I}}$, where $\rm\mathbb{I}$ is the identity operator on $\mathscr{H}$.
+We focus on the following specific class of POVMs.
+Definition 3.1.
+A quantum measurement is called rank-one if its measurement operators are rank-one, i.e., they take the outer-product form $Q_{i}=\mu\mu^{*}$ for some nonzero vectors $\mu\in\mathscr{H}$, where $\mu^{*}$ denotes the conjugate transpose of $\mu$.
+These associated vectors are referred to as measurement vectors.
+Remark 3.2.
+A rank-one POVM generalizes standard quantum measurement, as its measurement vectors $\{\mu_{j}\}$ need not to be normalized or orthogonal.
+There exists a fundamental connection between rank-one POVMs and the following so-called tight frames [4].
 Definition 3.3.
@@
 If $\alpha=1$, the tight frame is said to be normalized, otherwise, it is $\alpha$-scaled.
+More precisely, the connection between rank-one POVMs and tight frames is connected as below.
 Theorem 3.4.
@@
 In particular, the vectors $\varphi_{j}$ form a normalized tight frame for $\mathscr{H}$ if and only if they are the measurement vectors of a rank-one POVM on $\mathscr{H}$.
-When no confusion can arise, we will use the same notation for the POVMs and the corresponding set of measurement vectors.
-Definition 3.3 ($s$-order incompatibility for POVMs).
-Two rank-one POVMs $A$ and $B$ are called $s$-order incompatible if ${\rm spark}(A,B)=s$, where $(A,B)$ are the measurement vector matirx.
-For a pure state |$\ket{\varphi}$, let $\{\langle a_{j}|\varphi\rangle\}$ and $\{\langle b_{k}|\varphi\rangle\}$ denote the coefficients under the POVMs.
-Denote $n_{A}(\ket{\varphi})$ for the number of non-zero elements in $\{\langle a_{j}|\varphi\rangle\}$ and $n_{B}(\ket{\varphi})$ as the number of non-zero elements in $\{\langle b_{k}|\varphi\rangle\}$.
-Then the following sharp support uncertainty relation holds.
-Theorem 3.4 (support uncertainty relation).
-For $s$-order incompatible rank-one POVMs $A$ and $B$, Then for any non-zero pure state |$\ket{\varphi}$, we have
-$n_{A}(\ket{\varphi})+n_{B}(\ket{\varphi})\geq s.$
+Remark 3.5.
+For a pure state |$\ket{\varphi}$, the probability of observing the $i$th outcome is
+$p(i)=\langle\varphi,Q_{i}\varphi\rangle=|\langle\mu_{i},\varphi\rangle|^{2}.$
+obviously, the probabilities $p(i)$ sum to $1$.
+Using this connection, we classify the incompatibility of two rank-one POVMs in terms of tight frames.
+Definition 3.6.
+Suppose $A=\{a_{k},1\leq k\leq m\}$ and $B=\{b_{j},1\leq j\leq n\}$ are two tight frames in $\mathscr{H}$, with frame constant $\alpha$ and $\beta$, respectively.
+If an integer $s$ satisfies the following conditions:
+  1. (1)
+    For any non-empty subsets $S\subseteq I$ and $J\subseteq J$ with $|S|+|T|<s$, and for any non-zero $x\in\mathscr{H}$, at most one of the following holds,
+    $\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2}\quad{\rm and}\quad\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}.$
+  2. (2)
+    There exist non-empty subsets $S\subseteq I$ and $T\subseteq J$ with $|S|+|T|=s$, and a a nonzero $x\in\mathscr{H}$, such that
+    $\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2}\quad{\rm and}\quad\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}.$
+Then the frames $A$ and $B$ are said to be $s$-order incompatible.
+Remark 3.7.
+This notion is also linear algebraic. The two conditions in the definition are often more practical to be verified by considering the spanning set of frame vectors.
+Remark 3.8.
+When $A$ and $B$ are orthonormal bases, we recover the $s$-order incompatibility in Definition 2.1.
+Importantly, it should be noted that if $A$ and $B$ are $s$-order incompatible, then
+(3.1) $s\geq{\rm spark}(A,B).$
+This follows from the fact that frames are typically overcomplete. We provide an example.
+Example 3.9.
+Consider the tight frames $A=\{a_{1},a_{2}\}$ and $B=\{b_{1},b_{2},b_{3}\}$ (see Figure 1), where
+  1. (1)
+    $a_{1}=(1,0)$, $a_{2}=(0,1)$;
+  2. (2)
+    $b_{1}=\left(0,\frac{\sqrt{2}}{2}\right)$, $b_{2}=\left(\frac{\sqrt{2}}{2},\frac{1}{2}\right)$ and $b_{3}=\left(\frac{\sqrt{2}}{2},-\frac{1}{2}\right)$.
+(a) (1)
+(b) (2)
+Figure 1. Frames vectors of $A$ and $B$
+It is not hard to verify
+$\min(n_{A}(x)+n_{B}(x))=3,\quad{\rm whereas}\quad{\rm spark}(A,B)=2.$
+At the end of this section, we show the incompatibility and the support uncertainty relation are closely related as expected.
+For a pure state |$\ket{\varphi}$, let $n_{A}(\ket{\varphi})$ and $n_{B}(\ket{\varphi})$ denote the number of non-zero elements in the sequences $\{\langle a_{j},\varphi\rangle\}_{j=1}^{m}$ and $\{\langle b_{k},\varphi\rangle\}_{k=1}^{n}$, respectively.
+Theorem 3.10 (tight support uncertainty relation).
+Let $A$ and $B$ be two $s$-order incompatible tight frames. Then, for any non-zero pure state |$\ket{\varphi}$, we have
+$n_{A}(\ket{\varphi})+n_{B}(\ket{\varphi})\geq s,$
+where $s$ is optimal.
 Proof.
-It follows from the property of spark, as detailed in [11]. ∎
-This yields an equivalent definition of the $s$-order incompatibility for a pair of POVMs or for general tight frames.
-Definition 3.5.
-Suppose $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ are two tight frames in $\mathscr{H}$, with frame constant $\alpha$ and $\beta$. If the integer $s$ satisfies
+Suppose $\min_{\ket{\varphi}}\{n_{A}(\ket{\varphi})+n_{B}(\ket{\varphi})\}=t$. This means the following two:
   1. (1)
-    For any non-empty subsets $S\subseteq I$ and $J\subseteq J$, satisfying $|S|+|T|<s$, such that for any non-zero $x\in\mathscr{H}$,
-    $\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2},\quad{\rm and}\quad\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}$
-    do not hold simultaneously,
+    There exists a pure state |$\ket{\varphi}$, such that $n_{A}(\ket{\varphi})+n_{B}(\ket{\varphi})=t$.
+Namely for such |$\ket{\varphi}$, let $S=\{k:\langle a_{k},\varphi\rangle\neq 0\}$ and $T=\{j:\langle x,b_{j}\rangle\neq 0\}$, we have $|S|+|T|=t$,
+    $\sum_{k\in S}|\langle a_{k},\varphi\rangle|^{2}=\alpha\|x\|^{2}\quad{\rm and}\quad\sum_{j\in T}|\langle b_{j},\varphi\rangle|^{2}=\beta\|x\|^{2}$
+    according to the definition of tight frames.
   2. (2)
-    There exist non-empty subsets $S\subseteq I$ and $T\subseteq J$, satisfying $|S|+|T|=s$, such that there exists a non-zero $x\in\mathscr{H}$, the following
-    $\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2},\quad{\rm and}\quad\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}$
-    hold simultaneously.
-Then $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ are called $s$-order incompatible.
-4. Uncertainty relations from $s$-order incompatibility
-In this section, we establish new uncertainty relation based on the concept of s-order incompatibility.
-By leveraging the spark of measurement matrices, we derive lower bounds on the support of quantum states under different measurement settings, extending traditional uncertainty principles to a broader class of quantum measurements.
-Let $\{a_{k}\}_{k=1}^{m}$ and $\{b_{j}\}_{j=1}^{n}$ be two frames in a Hilbert space $\mathscr{H}$, where each set corresponds to a rank-one positive operator-valued measure (POVM).
+    There does not exist a nonzero state |$\ket{\phi}$ such that $n_{A}(\ket{\phi})+n_{B}(\ket{\phi})<t$.
+It is equivalent to say for any nonempty sets $S\subset I$ and $T\subset J$ satisfying $|S|+|T|<t$, and for any non-zero $x\in\mathscr{H}$, at most one of the following holds
+    $\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2},\quad{\rm and}\quad\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}.$
+Comparing with Definition 3.6, the above indeed means that $A$ and $B$ are $t$-order incompatible. Therefore $t=s$, the proof is complete. ∎
+4. Uncertainty relation from $s$-order incompatibility
+The main aim of this section is to derive an uncertainty relation based on $s$-order incompatibility.
+While these bounds can be obtained using the support uncertainty relation in Theorem 3.10 combined with a standard compactness argument, the constant derived through this approach remains unclear (as also noted in [5]).
+We study it from a new linear algebra perspective.
+Let $\{a_{k}\}_{k=1}^{m}$ and $\{b_{j}\}_{j=1}^{n}$ be two tight frames in a Hilbert space $\mathscr{H}$. After normalization, each set corresponds to a rank-one POVM.
 Definition 4.1 (Minimal Reconstruction Number).
@@
 $t_{\min}=\min\{t\,|\,\forall S\subsetneq I,T\subsetneq J,|S|+|T|>t,\operatorname{span}\left(\{a_{k}\}_{k\in S}\cup\{b_{j}\}_{j\in T}\right)=\mathscr{H}\}.$
-For any pair of $s$-order incompatible tight frames, the relationship between $s$ and $t_{\min}$ is given by the following theorem:
+It is shown that the quantities $s$ and $t_{\min}$ are intrinsically related.
 Theorem 4.2.
-If two tight frames $A$ and $B$ are $s$-order incompatible, then
+Suppose two tight frames $A$ and $B$ are $s$-order incompatible. Then
 $s+t_{\min}=m+n.$
@@
 Since the frames $\{a_{k}\}_{k=1}^{m}$ and $\{b_{j}\}_{j=1}^{n}$ are $s$-order incompatible, thus by the equivalent Definition 3.6, for any non-zero vector $x\in\mathscr{H}$, the following two identities
-$\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2},\quad\mbox{and}\quad\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}$
+$\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2}\quad\mbox{and}\quad\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}$
 cannot simultaneously hold. This implies that
@@
 $\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}=\alpha\|x\|^{2},\quad\mbox{and}\quad\sum_{j\in T}|\langle x,b_{j}\rangle|^{2}=\beta\|x\|^{2}.$
-This means that for any $S^{c}\subsetneq I$ and $T^{c}\subsetneq J$ satisfying $|S^{c}|+|T^{c}|=m+n-s$, there exists non-zero $x\in\mathscr{H}$ such that
+This implies that for any $S^{c}\subsetneq I$ and $T^{c}\subsetneq J$ satisfying $|S^{c}|+|T^{c}|=m+n-s$, there exists non-zero $x\in\mathscr{H}$ such that
 $\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}+\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}=0.$
-Thus $\left({\rm span}\left(\{a_{k}\}_{k\in S}\cup\{b_{j}\}_{j\in T}\right)\right)^{\perp}\neq\{0\}$, i.e.
-${\rm span}\left(\{a_{k}\}_{k\in S}\cup\{b_{j}\}_{j\in T}\right)\neq\mathscr{H}.$
+Thus $\left({\rm span}\left(\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}\right)\right)^{\perp}\neq\{0\}$, i.e.,
+${\rm span}\left(\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}\right)\neq\mathscr{H}.$
 Hence $t_{\min}\geq m+n-s$.
@@
 Remark 4.3.
-The quantity $t_{\min}$ establishes a fundamental limit on the number of measurements needed to reconstruct a quantum state from two rank-one POVMs.
+The quantity $t_{\min}$ provides a lower bound (i.e., $t_{\min}+1$) on the number of measurements required to reconstruct a quantum state from two rank-one POVMs.
+By Theorem 4.2, for any non-empty subsets $S\subset I$ and $T\subset J$ satisfying $|S|+|T|<s$, we have
+$\operatorname{span}\left(\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}\right)=\mathscr{H}.$
+Thus the set $\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}$ forms a frame in $\mathscr{H}$ (by [1, Lemma 1.2]), whose lower frame bound is denoted by $C_{S,T}$.
+Furthermore, denoting $C_{s}=\min_{|S|+|T|<s}C_{S,T}$.
+The following uncertainty relation follows.
 Theorem 4.4.
-Let $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ in $\mathscr{H}$ be two $s$-order incompatible tight frames. Then there exists a constant $C$ such that
+Let $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ be two $s$-order incompatible tight frames in $\mathscr{H}$.
+Then there exists a constant $C$, such that for all subsets $S\subseteq I$ and $T\subseteq J$ satisfying $|S|+|T|<s$, we have
 (4.1) $\|x\|^{2}\leq C\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}+\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right).$
-for all subsets $S\subseteq I$ and $T\subseteq J$ with $|S|+|T|<s$.
-Proof.
-By Theorem 4.2, for any non-empty subsets $S\subset I$ and $T\subset J$ satisfying $|S|+|T|<s$, we know that
-$\operatorname{span}\left(\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}\right)=\mathscr{H}.$
-Thus the set $\{a_{k}\}_{k\in S^{c}}\cup\{b_{j}\}_{j\in T^{c}}$ forms a frame for $\mathscr{H}$ (by [14, Lemma 1.2]), whose lower frame bound is denoting by $C_{S,T}$.
-Denoting $C_{1}=\min C_{S,T}$, the desired constant can be chosen as $C=1/\min\{\alpha,\beta,C_{1}\}$.
-∎
+where $C=1/\min\{\alpha,\beta,C_{s}\}$, in which $\alpha$ and $\beta$ are the frame constants.
 Remark 4.5.
-If the columns of the matrix $M$ are normalized to unit norm, then the lower bound of its spark can be given by its dictionary coherence
-${\rm spark}(M)\geq 1+\frac{1}{\mu(M)},$
-where the dictionary coherence $\mu(A)$ is defined as the maximum correlation between any to columns
-$\mu(M)=\max_{m\neq n}|\langle a_{m},a_{n}\rangle|.$
-It is seen that our uncertainty inequality provides a more essential interpretation for Ghobber-Jaming uncertainty relation in [3] in terms of the coherence $\mu(M)$.
-5. Ghobber-Jaming uncertainty relation for frames
+If the sets $S$ and $T$ are fixed, the constant in (4.1) can be chosen as $C=1/\min\{\alpha,\beta,C_{S,T}\}$.
+For a matrix $M$, the dictionary coherence $\mu(A)$ is defined as the maximum correlation between any two distinct columns
+(4.2) $\mu(M)=\max_{m\neq n}|\langle a_{m},a_{n}\rangle|.$
+If the columns of $M$ are normalized to unit norm, a lower bound of its spark is given in terms of its dictionary coherence by
+${\rm spark}(M)\geq 1+\frac{1}{\mu(M)}.$
+Thus, the uncertainty relation (4.1) is closely related to the Ghobber-Jaming inequality [5].
+However, by inequality (3.1), it is seen that (4.1) is more general, while the constant provided in [5] is explicit particularly for orthogonal bases.
+5. Ghobber-Jaming uncertainty for frames
+This section is dedicated to generalizing the Ghobber-Jaming uncertainty inequality in [5] for two frames $A=\{a_{k}\}_{k=1}^{m}$ and $B=\{b_{j}\}_{j=1}^{n}$ (not necessarily tight) in terms of their coherence, which is given by
+$M(A,B)=\max_{k,j}|\langle a_{k},b_{j}\rangle|$
+with explicit constant. This question was raised at the end of [13].
 Theorem 5.1.
-Let $\{a_{k},1\leq k\leq m\}$ and $\{b_{j},1\leq j\leq n\}$ be two frames in $\mathscr{H}$, with lower and upper frame bounds $\alpha_{1},\beta_{1}$ and $\alpha_{2},\beta_{2}$, respectively.
-Suppose that for any subsets $S\subseteq\{1,\dots,m\}$ and $T\subseteq\{1,\dots,n\}$, satisfying $S\subseteq T$, the condition
-$\frac{\mu^{2}}{\alpha_{1}}<\frac{\beta_{2}}{T}$
-holds, where $\mu(M)=\max_{m\neq n}|\langle a_{m},a_{n}\rangle|$. Then for all $x\in H$,
-$C(S,T)\|x\|\leq\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}.$
+Let $A$ and $B$ be two frames in $\mathscr{H}$, with lower and upper frame bounds $\alpha_{1},\beta_{1}$ for $A$ and $\alpha_{2},\beta_{2}$ for $B$, respectively.
+Suppose that for any subsets $S\subseteq\{1,\dots,m\}$ and $T\subseteq\{1,\dots,n\}$, satisfying
+$|S||T|<\frac{\beta_{1}}{\alpha_{2}}\cdot\frac{1}{M(A^{*},B)^{2}}.$
+Then, for any $x\in\mathscr{H}$, we have
+$C(S,T)\|x\|\leq\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}},$
+where
+$\begin{split}C(S,T)=&\left(1-\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}M(A^{*},B)|S|^{\frac{1}{2}}|T|^{\frac{1}{2}}\right)\\ &\times\left(\max\left\{\beta_{1}^{-1/2},\left(1+\left(\frac{\beta_{2}}{\beta_{1}}\right)^{\frac{1}{2}}\right)\alpha_{1}^{-\frac{1}{2}}mM(A^{*},A)\right\}\right)^{-1}.\end{split}$
 Proof.
 Let $x=x_{1}+x_{2}$, where
-$x_{1}=\sum_{k\in S}\langle x,a_{k}\rangle a_{k}^{*},\quad x_{2}=\sum_{k\in S^{c}}\langle x,a_{k}\rangle a_{k}^{*},$
-in which $a_{k}^{*}$ is the canonical dual frame.
-$\begin{split}\sum_{j\in T}|\langle x_{1},b_{j}\rangle|^{2}=\sum_{j\in T}\left|\sum_{k\in S}\langle x,a_{k}\rangle\langle a_{k}^{*},b_{j}\rangle\right|^{2}\\ \leq\sum_{j\in T}\left(\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}\right)\left(\sum_{k\in S}\left|\langle a_{k}^{*},b_{j}\rangle\right|\right)\\ \leq\mu^{2}\alpha_{2}|T|\|x\|^{2}.\end{split}$
-Thus,
-$\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}=\sum_{j=1}^{n}\left|\langle x_{1},b_{j}\rangle\right|^{2}-\sum_{j\in T}|\langle x_{1},b_{j}\rangle|^{2}\geq\beta_{1}\|x_{1}\|^{2}-\alpha_{2}\mu^{2}\mu^{2}|T|\|x\|^{2}$
-This implies
-$\begin{split}\beta_{1}^{\frac{1}{2}}\|x_{1}\|&\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}+\alpha_{2}\mu^{2}|T|\|x\|^{2}\right)^{\frac{1}{2}}\\ &\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\alpha_{2}^{\frac{1}{2}}\mu|T|^{\frac{1}{2}}\|x\|\\ &\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x_{2},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\alpha_{2}^{\frac{1}{2}}\mu|T|^{\frac{1}{2}}\|x\|\\ &\leq\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\beta_{2}^{\frac{1}{2}}\|x_{2}\|+\alpha_{2}^{\frac{1}{2}}\mu|T|^{\frac{1}{2}}\|x\|.\end{split}$
+$x_{1}=\sum_{k\in S}\langle x,a_{k}\rangle a_{k}^{*},\quad{\rm and}\quad x_{2}=\sum_{k\in S^{c}}\langle x,a_{k}\rangle a_{k}^{*},$
+in which $a_{k}^{*}$ is the canonical dual frame. For $x_{1}$, we have
+$\begin{split}\sum_{j\in T}|\langle x_{1},b_{j}\rangle|^{2}&=\sum_{j\in T}\left|\sum_{k\in S}\langle x,a_{k}\rangle\langle a_{k}^{*},b_{j}\rangle\right|^{2}\\ &\leq\sum_{j\in T}\left(\sum_{k\in S}|\langle x,a_{k}\rangle|^{2}\right)\left(\sum_{k\in S}\left|\langle a_{k}^{*},b_{j}\rangle\right|^{2}\right)\\ &\leq\alpha_{2}M(A^{*},B)^{2}|S||T|\|x\|^{2}.\end{split}$
+Therefore,
+$\begin{split}\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}&=\sum_{j=1}^{n}\left|\langle x_{1},b_{j}\rangle\right|^{2}-\sum_{j\in T}|\langle x_{1},b_{j}\rangle|^{2}\\ &\geq\beta_{1}\|x_{1}\|^{2}-\alpha_{2}M(A^{*},B)^{2}|S||T|\|x\|^{2}.\end{split}$
+It yields that
+(5.1) $\begin{split}\beta_{1}^{\frac{1}{2}}\|x_{1}\|\leq&\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}+\alpha_{2}M(A^{*},B)^{2}|S||T|\|x\|^{2}\right)^{\frac{1}{2}}\\ \leq&\left(\sum_{j\in T^{c}}|\langle x_{1},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\alpha_{2}^{\frac{1}{2}}M(A^{*},B)|S|^{\frac{1}{2}}|T|^{\frac{1}{2}}\|x\|\\ \leq&\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x_{2},b_{j}\rangle|^{2}\right)^{\frac{1}{2}}\\ &+\alpha_{2}^{\frac{1}{2}}M(A^{*},B)|S|^{\frac{1}{2}}|T|^{\frac{1}{2}}\|x\|\\ \leq&\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\beta_{2}^{\frac{1}{2}}\|x_{2}\|+\alpha_{2}^{\frac{1}{2}}M(A^{*},B)|S|^{\frac{1}{2}}|T|^{\frac{1}{2}}\|x\|.\end{split}$
 On the other hand,
-$\begin{split}\alpha_{1}\|x_{2}\|^{2}&\leq\sum_{k=1}^{m}|\langle x_{2},a_{k}\rangle|^{2}\leq\sum_{k=1}^{m}\left(\sum_{i\in S^{c}}|\langle x,a_{i}\rangle|^{2}\right)\left(\sum_{i\in S^{c}}|\langle a_{i}^{*},a_{k}\rangle|^{2}\right)\\ \leq m^{2}M(A^{*},A)^{2}\sum_{k\in s^{c}}|\langle x,a_{k}\rangle|^{2}\end{split}$
-Thus we have
-$\begin{split}\|x\|\leq\|x_{1}\|+\|x_{2}\|&\leq\frac{1}{\beta_{1}^{\frac{1}{2}}}\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\left(1+\left(\frac{\beta_{2}}{\beta_{1}}\right)^{1/2}\right)\|x_{2}\|+\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}\mu|T|^{1/2}\|x\|\end{split}$
-we obtain
-$\begin{split}C(S,T)\|x\|\leq\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{1/2}\end{split}$
+(5.2) $\begin{split}\alpha_{1}\|x_{2}\|^{2}&\leq\sum_{k=1}^{m}|\langle x_{2},a_{k}\rangle|^{2}\leq\sum_{k=1}^{m}\left(\sum_{i\in S^{c}}|\langle x,a_{i}\rangle|^{2}\right)\left(\sum_{i\in S^{c}}|\langle a_{i}^{*},a_{k}\rangle|^{2}\right)\\ &\leq m^{2}M(A^{*},A)^{2}\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}.\end{split}$
+By (5.1) and (5.2), we have
+$\begin{split}\|x\|\leq\|x_{1}\|+\|x_{2}\|\leq&\frac{1}{\beta_{1}^{\frac{1}{2}}}\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{\frac{1}{2}}+\left(1+\left(\frac{\beta_{2}}{\beta_{1}}\right)^{1/2}\right)\|x_{2}\|\\ &+\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}M(A^{*},B)|S|^{\frac{1}{2}}|T|^{\frac{1}{2}}\|x\|.\end{split}$
+Regrouping terms, we obtain
+$\begin{split}C(S,T)\|x\|\leq\left(\sum_{k\in S^{c}}|\langle x,a_{k}\rangle|^{2}\right)^{\frac{1}{2}}+\left(\sum_{j\in T^{c}}|\langle x,b_{j}\rangle|^{2}\right)^{1/2},\end{split}$
 where
-$C(S,T)=\left(1-\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}\mu|T|^{1/2}\right)\left(\max\{\frac{1}{\beta_{1}^{1/2}},\left\{\beta_{1}^{-\frac{1}{2}},\left((\frac{\beta_{2}}{\alpha_{1}\beta_{1}})^{\frac{1}{2}}+\alpha_{1}^{-\frac{1}{2}}\right)mM(A^{*},A)\right\}\right)^{-1}.$
-6. Incompatibility of multiple bases
-We consider the incompatibility of multiple bases in this section.
+$\begin{split}C(S,T)=&\left(1-\left(\frac{\alpha_{2}}{\beta_{1}}\right)^{1/2}M(A^{*},B)|S|^{\frac{1}{2}}|T|^{\frac{1}{2}}\right)\\ &\times\left(\max\left\{\beta_{1}^{-1/2},\left(1+\left(\frac{\beta_{2}}{\beta_{1}}\right)^{\frac{1}{2}}\right)\alpha_{1}^{-\frac{1}{2}}mM(A^{*},A)\right\}\right)^{-1}.\end{split}$
+Remark 5.2.
+It is interesting to further improve the bound in terms of
+$\mu_{r}(A,B)=\max_{1\leq j\leq n}\left(\sum_{k=1}^{m}|\langle a_{k},b_{j}\rangle|^{r^{\prime}}\right)^{\frac{r}{r^{\prime}}},$
+where $r\in[1,2]$ and $11/r+1/r^{\prime}=$ 1, similarly as done in [15].
+6. Incompatibility of multiple tight frames
+The main goal of this section is to extend the notion of $s$-order incompatibility for multiple tight frames ($n\geq 3$).
+When dealing with multiple frames, a suitable approach should consider the overall coherence of the entire collection, rather than analyzing each pair separately and subsequently combining the results.
+To illustrate the differences, we conclude this section with an example.
 Definition 6.1.
-Orthonormal bases $A_{k}=\{\ket{a^{(k)}_{j}}\}_{j=1}^{n}$ are $s$-order incompatible if:
+Suppose $A_{i}=\{a_{k}^{(i)},1\leq k\leq m_{i}\}$ are $n$ tight frames. They are said to be $s$-order incompatible if:
   1. (1)
-    For any nonempty $S_{k}\subset A_{k}$ with $\sum_{k=1}^{n}|S_{k}|<s$, $\bigcap_{k=1}^{n}{\rm span}\{S_{k}\}=\{0\}.$
+    For any non-empty subsets $S_{i}\subset I_{i}$ satisfying $\sum_{i=1}^{n}|S_{i}|<s$, and for any non-zero $x\in\mathscr{H}$, the following
+    $\sum_{k\in S_{i}}\left|\left\langle x,a_{k}^{(i)}\right\rangle\right|^{2}=\alpha_{i}\|x\|^{2},$
+    cannot hold for all $i=1,2,\ldots,n$.
   2. (2)
-    There exist subsets $S_{k}\subset A_{k}$, such that $\sum_{k=1}^{n}|S_{k}|=s$ and $\bigcap_{k=1}^{n}{\rm span}\{S_{k}\}\neq\{0\}.$
+    There exist non-empty subsets $S_{i}\subseteq I_{i}$ with $\sum_{i=1}^{n}|S_{i}|=s$, and a non-zero $x\in\mathscr{H}$, such that
+    $\sum_{k\in S_{i}}\left|\left\langle x,a_{k}^{(i)}\right\rangle\right|^{2}=\alpha_{i}\|x\|^{2}$
+    holds for all $i=1,2,\ldots,n$.
 Theorem 6.2.
-For $s$-order incompatible bases $\{A_{k}\}_{k=1}^{n}$, any non-zero pure state |$\ket{\varphi}$ if and only if
+The tight frames $A_{1},A_{2},\cdots,A_{n}$ are $s$-order incompatible if and only if for any non-zero pure state |$\ket{\varphi}$, it holds that
 $\sum_{k=1}^{n}n_{A_{k}}(\ket{\varphi})\geq s.$
+The constant $s$ here is optimal.
 Proof.
-Let
-$n_{\min}=\min_{\ket{\varphi}\neq 0}\sum_{k=1}^{n}n_{A_{k}}(\ket{\varphi}).$
-Then if $n_{\min}=s$, there exists a pure state |$\ket{\psi}$ such that
-$\sum_{k=1}^{n}n_{A_{k}}(\ket{\psi})=s,$
-and there does not exist a pure state |$\ket{\psi^{\prime}}$ such that $\sum_{k=1}^{n}n_{A_{k}}(\ket{\psi^{\prime}})<s$.
+Denote
+$n_{\min}=\min_{\ket{\varphi}\neq 0}\sum_{k=1}^{n}n_{A_{k}}(\ket{\varphi})=t.$
+Then there exists a pure state |$\ket{\psi}$ such that
+$\sum_{k=1}^{n}n_{A_{k}}(\ket{\psi})=t,$
+and there does not exist a pure state |$\ket{\psi^{\prime}}$ such that $\sum_{k=1}^{n}n_{A_{k}}(\ket{\psi^{\prime}})<t$.
 For such |$\ket{\psi}$, there exist nonempty subsets $S_{A_{k}}\subseteq A_{k}$ such that $|S_{A_{k}}|=n_{A_{k}}(\ket{\psi})$, and |$\ket{\psi}\in\bigcap_{k=1}^{n}{\rm span}(S_{A_{k}})$.
 The nonexistence of such |$\ket{\psi^{\prime}}$ implies that there do not exist nonempty subsets $S_{A_{k}}\subseteq A_{k}$ such that $|S_{A_{k}}|=n_{A_{k}}(\ket{\psi^{\prime}})$, and |$\ket{\psi^{\prime}}\in\bigcap_{k=1}^{n}{\rm span}(S_{A_{k}})$.
-These two conditions precisely coincide with (1.1) and (1.2) in Definition 1.
+These two conditions precisely coincide with the two conditions in Definition 6.1, which implies that $t=s$.
 Then the claim follows.
 $\square$
+We show that the multiple cases ($n\geq 3$) can not be derived from the classical two-frame case.
+Theorem 6.3.
+Suppose tight frames $A_{1},A_{2},\cdots,A_{n}$ are $s$-order incompatible and each pair $A_{i}$ and $A_{j}$ are $s_{ij}$-order incompatible, then we have
+(6.1) $\frac{1}{2}\sum_{i<j}s_{ij}\leq s.$
+Proof.
+Since $A_{1},A_{2},\cdots,A_{n}$ are $s$-order incompatible, then there exists a nonzero $x\in\mathscr{H}$ such that
+(6.2) $n_{A_{1}}(x)+n_{A_{2}}(x)+\ldots+n_{A_{n}}(x)\geq s.$
+On the other side, since $A_{i}$ and $A_{j}$ are $s_{ij}$-order incompatible, we have
+(6.3) $n_{A_{i}}(x)+n_{A_{j}}(x)\geq s_{ij}.$
+Combining (6.2) and (6.3), we get the result.
+The following example shows that the strict inequality in (6.1) can occur.
+Example 6.4.
+Consider the frames $A=\{a_{1},a_{2},a_{3}\}$, $B=\{b_{1},b_{2},\ldots,b_{5}\}$ and $C=\{c_{1},c_{2},c_{3},c_{4}\}$, where
+  1. (1)
+    $a_{1}=(2\sqrt{3},1),a_{2}=(-2,-\sqrt{3}),a_{3}=(0,2\sqrt{3});$
+  2. (2)
+    $b_{1}=(1,1),b_{2}=(\frac{1}{2},\frac{1}{2}),b_{3}=(0,1),b_{4}=(2,-\sqrt{3}),$
+    $b_{5}=(-\left(2\sqrt{3}-\frac{5}{4}\right)^{\frac{1}{2}},-\left(2\sqrt{3}-\frac{5}{4})^{\frac{1}{2}}\right)$;
+  3. (3)
+    $c_{1}=(2,-\sqrt{3}),c_{2}=(2\cdot 3^{\frac{1}{4}},2\cdot 3^{\frac{1}{4}}),c_{3}=(-2,\sqrt{3}),c_{4}=(0,-\sqrt{2})$.
+(a) (1)
+(b) (2)
+(c) (3)
+Figure 2. frames vectors of $A,B$ and $C$
+Straightforward computations show that $s_{AB}=5$, $s_{BC}=5$, $s_{AC}=4$ and $s=8$. Therefore,
+$\frac{s_{AB}+s_{BC}+s_{AC}}{2}<s.$
+This means that the cases for multiple measurements $n\geq 3$ differs from the uncertainty relation for two bases.
 7. Conclusion
-By equating $s$-order incompatibility with the spark, we unify quantum measurement theory with compressed sensing. Key contributions include:
-  • POVM Incompatibility: Generalized to non-projective measurements via tight frames.
-  • Sharp Uncertainty Bounds: Derived from the spark, improving coherence-based estimates.
-  • Multi-Basis Framework: Extended support uncertainty to multi-observable systems.
-Future directions include applications in quantum tomography, cryptographic protocol design, and infinite-dimensional spark generalizations.
+In this paper, we show that the $s$-order incompatibility of two orthonormal bases coincides with the spark of matrix formed by the measurement vectors, thereby linking quantum measurement with compressed sensing.
+The main contribution of this work is the classification of the incompatibility of two and multiple ($n\geq 3$) tight frames (rank-one POVMs), leading to sharp support uncertainty relations and new generalizations of Ghobber-Jaming uncertainty relation.
+Potential directions for future research include applications in quantum tomography, cryptographic protocol design, and explicit calculations of the order of incompatibility.
+Acknowledgments
+This work was partially supported by NSFC Grant No.12101451.